# Impact of evanescent scattering modes and finite dispersion in superconducting junctions

### v1

@@ VI.1 Normal metal scattering with evanescent modes
At this stage, we can explicitly evaluate all scattering amplitudes for the L-junction problem. In particular, this allows us to numerically check relation (80), which followed from the analytical treatment presented in Sec. V, assuming ${\Delta/\mu_{2}\ll 1}$ while still allowing for ${\mu_{2}\to 0}$. As established in the same section, if this constraint on the scattering matrix holds, we can deduce that energy solutions at ${E=\Delta}$ exist in the ABS spectrum at ${\phi=0}$  notably, without requiring anything beyond solving the scattering matrix itself. To this end, we compute the deformation term ${\Omega=\partial S_{e}^{\text{pp}}(S_{e}^{\text{pp}})^{*}-S_{e}^{\text{pp}}(\partial S_{e}^{\text{pp}})^{*}}$, the compensation term ${\Sigma^{\ast}/\left(4\mu_{2}\right)}$ and the difference of the two ${\delta=\Omega/\mathrm{i}-\mathrm{i}\Sigma^{\ast}/\left(4\mu_{2}\right)}$. If $\delta=0$, then the constraint of Eq. (80) would hold exactly. As shown in Table 1, the deviation $\delta$ is small but finite and the relative deviation ${\left\|\delta\right\|/\left\|\Omega\right\|}$ indeed converges with increasing finite dispersion $\mu_{2}\to 0$, readily confirming relation (80). It is interesting to note that the absolute deviation $\delta$ stays constant. This can be attributed to the finite size of the central scattering region in the L-junction model (c.f. Figure 7), which obviously stays constant when changing the chemical potential. Consequently, a small residual probability density indeed remains within the finite scattering region. Importantly, these two contributions can be distinguished due to their respective behavior with respect to $\mu_{2}$ ($\sim 1/\mu_{2}$ versus constant). Thus, the here presented numerical analysis explicitly confirms one of our main results, Eq. (80), indicating the existence of surprising constraints on the energy-dependent scattering matrix involving evanescent modes, which do not follow from probability conservation. Below, we will go even further in the following sense. While Eq. (80) follows from perturbative arguments when $\Delta$ is sufficiently small, we will show below that the gap closing at $E=\Delta$ persists even non-perturbatively.
Therefore, in what follows we proceed towards a full calculation of the Andreev bound state spectrum. Here, the scattering matrix would in principle need to be evaluated for each energy $E$. However, we can significantly reduce computation time by exploiting the square root behavior deduced in Sec. III.4. Thus, the scattering problem only needs to be solved for the energies ${E=0}$ and ${E=\mu_{\text{crit}}}$, and the region in between, then, can be interpolated by the square root law. Consequently, the coefficients of the electron scattering matrix can be inferred as a function of energy by the following analytic relation
@@
explicitly capture the deviation across the whole energy range from ${E=0}$ to $\mu_{\text{crit}}$. This holds in complete analogy for ${\mathrm{Arg}\left[S^{i,j}_{\mathrm{e}}\left[E\right]\right]}$, such that the full electron scattering matrix can be assembled as ${S^{i,j}_{\mathrm{e}}\left[E\right]=\left|S^{i,j}_{\mathrm{e}}\left[E\right]\right|\cdot\mathrm{exp}\left[\mathrm{Arg}\left[S^{i,j}_{\mathrm{e}}\left[E\right]\right]\right]}$. The hole components of the scattering matrix can be calculated by using particle-hole symmetry. In Fig. 8, we explicitly illustrate the very good agreement between the square root behavior and the numerical calculation, with the example of the planar-planar reflection coefficient $R_{1,1}$. Here, the regime is such that the first channel is planar $\mu_{1}/\Delta\ll 0$, the second channel is evanescent with finite dispersion $\mu_{2}/\Delta\gtrapprox 0$, and the width of the arms is $N=5$ lattice sites.
VI.2 Andreev processes in the lattice model
The reflection processes at the SN-interfaces can, in the discrete model, be solved by a projection procedure similar to that of the normal scattering. However, since reflection at the interface is assumed to be translational invariant, as discussed in Section II, the problem can be separated such that transversal components drop out. Projecting onto the SN-interfaces of each contact at ${j_{\mu}=j_{\alpha}}$, see also Fig. 7, yields the following set of $2N$ conditions
which correspond to the differentiability constraints in the continuum limit. Depending on which side of the interface is involved, either the ansatz for the superconductor $|\psi^{\mathrm{SC}}_{n}\rangle$ or the normal conductor region $|\psi_{n}^{\mathrm{NC}}\rangle$ is applied, accordingly, the details of which are outlined in Appendix B. With the SN-interface being situated a finite distance of lattice points $j_{\alpha}$ from the scattering region, the wave function must be propagated in between by an operator $P$. In the absence of impurities (our default assumption for simplicity), this propagator $P$ consists of a diagonal matrix with elements ${P_{n,n}\left[j\right]=\mathrm{e}^{-\kappa_{\nu,n}j}}$, where the momentum ${\kappa_{\nu,n}}$ is determined as defined by Eq. (65). For instance, if $\psi^{\mathrm{in}}$ is the incident wave vector right at ${j=0}$, it would be subject to the eigenvector equation $PRP^{-1}S\psi^{\mathrm{in}}=\psi^{\mathrm{in}}$. In the standard Beenakker formalism as well as in our generalized framework above, the propagator is usually absorbed into the $S$-matrix (or, alternatively, in $R$).
While it is possible to derive a closed expression for the coefficients of the discrete SN interface, the respective formulas are impractically bulky (which is why they are not explicitly derived in Appendix B, and instead the lattice model coefficients are determined numerically). For comparison and in order to simplify the calculation, we can, however, use the analytic Andreev reflection- and backscattering coefficients $\alpha_{\nu,n}$, $\beta_{\nu,n}$ from the continuous model presented in Sec. IV, c.f. Eqs. (62) and (63), and simply insert the respective dispersion relations for electrons and holes from the lattice. It turns out that even if taking a conductor arm width of just $N=5$ lattice points the analytic approximation (solid curves) and the numerical computation (dashed curves) are in very good agreement (see Fig. 9). This is due to the fact that the chemical potential is chosen close to a change of the channel number, where the dispersion relation of the relevant channel (here the second channel, $n=2$) is to a very good accuracy parabolic. This fact allows us (in the subsequent section) to provide a much faster computation of the ABS spectrum, where the $S$-matrix can be interpolated by the square root law, Eq. (90), as discussed above, and using the analytic $R$-matrix expression from the continuous model. As we will show, the ABS spectra obtained with these two methods align to very high accuracy. In the following, we distinguish these two calculations by calling the former “numerical” (in that for each value of $E$, both $S$ and $R$ are computed numerically from the corresponding lattice boundary conditions) and the latter “analytic” (since $S$ and $R$ follow from a closed analytic expression).
One final important remark: while the regime considered here is close to the transition (with $\mu_{2}$ small) we no longer limit ourselves to $\Delta\ll\mu_{2}$. Consequently, and contrary to the perturbative results in Eqs. (70) and (71), the behavior of the evanescent $\alpha$ and $\beta$ coefficients is here much more complex ($\beta$ can no longer be neglected, and both coefficients depend strongly on $E$, see also Fig. 9).
VI.3 Non-perturbartive gap closing and long ballistic junctions
Figure 11: Impact of strongly dispersive evanescent cross-channel scattering on the bound states spectrum. The ABS spectra are computed in the extreme short junction limit ($j_{\alpha}=0$) with finite dispersion in the first evanescent channel (${\mu_{2}/\Delta=1.01}$, ${\mu_{1}/\Delta=-72.20}$) and $N=5$ transversal lattice sites. The black dotted curve depicts the standard result with energy dependent scattering and evanescent modes neglected, see Eq. (21). Panel a): if energy dependence (stemming from evanescent finite dispersion) is fully included in $S$, but evanescent channels are still discarded for $S$ and $R$, the spectra detach from the continuum with very good consistency between the numerical (dashed blue curve) and analytic (solid blue curve) computations. Panel b): if the calculation fully includes evanescent channels explicitly, (both in the scattering matrix $S$ as well as in the reflection matrix $R$) the detachment is compensated. Again, the numerical (dashed purple curve) and analytic (solid purple curve) treatment are in very good alignment.
Table 1: Numerical test of the validity of the $S$-matrix constraint (including evanescent modes) given in Eq. (80). We consider the case where the second channel is evanescent (${\mu_{2}>0}$) and the number of transversal lattice points is ${N=5}$. As $\mu_{2}/t$ decreases, dispersive effects become more and more pronounced. The table shows the terms ${\Sigma^{\ast}/\left(4\mu_{2}\right)}$ and ${\Omega=\partial S_{e}^{\text{pp}}(S_{e}^{\text{pp}})^{*}-S_{e}^{\text{pp}}(\partial S_{e}^{\text{pp}})^{*}}$ in Eq. (80) and its absolute deviation ${\delta=\Omega/\mathrm{i}-\mathrm{i}\Sigma^{\ast}/\left(4\mu_{2}\right)}$ in relation to the inverse energy scale $1/t$. ${\left\|\delta\right\|/\left\|\Omega\right\|}$ is the relative deviation with the matrix norm squared ${\left\|\delta\right\|^{2}=\sum_{n,m}\left|\delta_{n,m}\right|^{2}}$. The absolute deviation $\delta$ due to the finite extension of the scattering region stays constant, whereas the relative deviation ${\left\|\delta\right\|/\left\|\Omega\right\|}$ indeed decreases with increasing dispersion. We recall that the full scattering matrix including evanescent modes is evaluated at ${E=0}$, consistent with the discussion after Eq. (77).
As shown in Sec. VI.1, see in particular Tab. 1, the scattering matrix satisfies the analytical constraint, Eq. (80), derived in Sec. V, valid in the perturbative limit of $\Delta/\mu_{2}\ll 1$ (while $\mu_{2}\ll\left|\mu_{1}\right|$). We could therefore conclude already without explicit calculation of the ABS spectrum, that the gap closes at $E=\Delta$ in this perturbative limit. Here, we show by calculating the full ABS spectrum, that, notably, the effect of the gap closing persists even beyond the above perturbative limit. In fact, the only parameter that can still have an influence on whether a gap closing at $\phi=0$ is present or not is the distance between the scattering region and the SN interface, parametrized by the lattice position $j_{\alpha}$, see Fig. 7. Again, we assume for simplicity that there are no impurities within the conductor arms, such that the system for large $j_{\alpha}$ corresponds to a long ballistic junction.
Computing the normal region wave function, we can track the behavior of the evanescent component in between the scattering center and the SN-interface for different arm lengths $j_{\alpha}$. If $\psi^{\mathrm{in}}$ ($\psi^{\mathrm{out}}$) is the incoming (outgoing) wave vector right at the scattering region ${j=0}$, the first evanescent wave function component ${n=n_{\mathrm{P}}+1}$ then consists of the superposition ${\psi_{\nu,n}=\psi^{\mathrm{in}}_{\nu,n}P_{n,n}+\psi^{\mathrm{out}}_{\nu,n}P^{-1}_{n,n}}$, where $P$ is the propagator as introduced after Eq. (92). As Fig. 10 depicts, for increasing ${j_{\alpha}}$, real exponentials exiting the scattering region indeed increasingly drop off before reaching the SN-interface. Moreover, the qualitative behaviour of the wave function strongly depends on the ratio $\Delta/\mu_{2}$. In the nonperturbative regime $\Delta\approx\mu_{2}$ [Fig. 10a)], there is a much more pronounced diverging component within the evanescent mode as compared to the perturbative regime $\Delta\ll\mu_{2}$ [Fig. 10b)], resulting, accordingly, in a much slower decay of the wave function away from the scattering region.
If we apply the standard literature framework outlined in Sec. II, i.e., the scattering matrix is approximated for constant energy $S[E=0]$ and evanescent modes are discarded (such that the matrix $R$ is computed as in Eq. (11)), we arrive at the dotted black curve in Fig. 11. Here, the energy spectrum touches the gap $\Delta$ at $\phi=0$ simply due to the unitarity of the scattering matrix. Let us now include the full energy dependence of the scattering matrix $S\left[E\right]$, while still only retaining planar modes in the computation of the ABS spectrum. In essence, the finite dispersion effects of the evanescent channels are here only indirectly included in the computation of the scattering matrix, resulting in a strongly energy-dependent $S$ when approaching small values for $\mu_{2}$. Consequently, the bound states spectrum indeed detaches from the continuum at ${E=\Delta}$ (c.f. blue dashed curve in Figure 11a)), even though the electrons spend very little time within the scattering region. This detaching is in alignment with the chiral symmetry breaking argument brought forth in Sec. III.3. However, as amply explained throughout this work, this detaching is here not the correct result, due to the discarding of the evanescent modes in the boundary conditions for the ABS spectrum. Thus, the third and last step consists of fully taking into account the evanescent modes in both $S$ and $R$. To repeat, in alignment with the discussion after Eqs. (70) and (71), it is perfectly sufficient to only include the first evanescent mode, since all higher modes have vanishing dispersion (In fact, computing the ABS spectrum with two or three evanescent modes included, yields the exact same curves.). This full calculation indeed provides a bound states spectrum where the detaching is fully compensated, as depicted by the purple dashed curve in Fig. 11b). Again, we stress that here the revival of the gap closing at energies close to $\Delta$ is shown in the non-perturbative regime of $\Delta\approx\mu_{2}$ (and thus far beyond the regime considered in Sec. V). This result therefore very likely indicates the existence of non-perturbative constraints on the scattering matrix, beyond Eq. (80), allowing to relate scattering matrices (including evanescent scattering processes) at different energies, a question which can be pursued in follow-up works.
Figure 12: ABS spectra in the long junction limit ($j_{\alpha}=-300$) with ${N=5}$ transversal lattice points and finite dispersion in the first evanescent channel (${\mu_{2}/\Delta=1.01}$, ${\mu_{1}/\Delta=-72.20}$). The black dotted curve depicts the standard result with energy dependent scattering and evanescent modes neglected, see Eq. (21). The green dashed (solid) line depicts a numerical (analytic) calculation with only planar modes considered, but energy dependence fully included. The purple dotted curve is a numerical calculation, additionally taking evanescent modes into account. Consequently, in the long junction limit, the only relevant effect to take into account is the energy-dependence of $S$ due to strong dispersion. Contrary to the short junction limit (Fig. 11), here, the inclusion of evanescent modes in the $S$- and $R$-matrices does not preserve a gap closing.
Having determined the bound states spectra by exact numerical computation, we now compare to the aforementioned analytic approximations. To this end, the $R$-matrix is computed by the analytic expressions for the reflection- and backscattering coefficients $\alpha$$\beta$, Eqs. (62) and (63), with applying the discrete dispersion as explained in the previous section. Next, the scattering matrix ${S\left[E\right]}$ can be interpolated by the square root law, Eq. (90), along the energy range ${E\in\left[-\Delta,\Delta\right]}$. Solving for the ABS spectrum, the results (solid curves in Figs. 11 and 12) are in extremely good agreement with the numerical calculations (dashed curves) across all regimes. This indicates the validity of a number of helpful approximation schemes that allow for a more efficient computation of the ABS spectra in the here considered more general regimes, including evanescent modes, cross-channel scattering, and finite dispersive effects.
VII Conclusion and outlook
We provided a framework for the description of superconducting weak links in terms of a generic, energy-dependent scattering matrix, including evanescent modes and cross-channel coupling for general multichannel contacts. We provided an in depth analysis of the different origins of an energy-dependence of the scattering matrix, distinguishing in particular between a finite size of the scattering region itself (where the energy dependence can be linked to the inverse dwelling time of electrons) and a finite dispersion in the conductor arms. In particular, we argued that the latter mechanism can be (and in general is) large, when the chemical potential is tuned close to a regime where the number of planar channels changes. We further demonstrated that if energy-dependence in the scattering matrix is the only feature included in Beenakker’s determinant equation, then it predicts a detaching of the ABS spectrum between bound and extended quasiparticle states (i.e., at energies equal to $\Delta$). We further show that if the energy-dependence comes dominantly from a finite dispersion (while the actual dwelling time remains neglible), then the detaching is spurious – and a more complete theoretical treatment still predicts a gap closing. By means of an effective Hamiltonian description of the weak link, we provide a connection between the gapping and the breaking of a chiral symmetry. Depending on whether the dominant dispersive channel is planar or evanescent, there are distinctly different lines of argumentation as to how the chiral symmetry, and thus the gap closing, are restored. For a dispersive planar mode, we argue that there are probability conservation constraints on the scattering matrix beyond unitarity, which can be derived through time-dependent considerations of the scattering process. Crucially, we show that probability conservation arguments are not applicable if the dispersion stems from an evanescent mode, since not all scattering wave functions containing those modes can be normalized. Nonetheless, by careful consideration of the generalized formalism to compute the ABS eigenspectrum, we find new types of constraints on the energy-dependent scattering matrix including coupling to evanescent modes, valid in the perturbative limit of weak dispersion compared to $\Delta$. Finally, this constraint, and the gap closing in a nonperturbative regime are explicitly demonstrated numerically, with the example of a ballistic L-junction. We find that the emergence of gaps in the ABS spectrum at zero phase bias ($\phi-0$) are only physical when increasing the length of the conductor arms, such that evanescent modes decay before they collide with the SN interface.
Finally, we expect this work to be of relevance for multi-pronged future research efforts. On the one hand, energy gaps, especially those close to the continuum, play an important role in the proper description of time-dependently driven junctions, either regarding Landau-Zener transitions coupling the discrete and continuous states (similar in spirit to Ref. [25]) or with respect to their dissipative behavior [8, 10]. On the other hand, unitarity and other constraints of the scattering matrix are well-known to play a crucial role in the context of random matrix theory [57, 58, 59, 60, 17, 21]. This work is to the best of our knowledge the first to consider constraints regarding the energy-dependence of the scattering matrix. We therefore consider it likely that our work provides a basis for subsequent studies linking the scattering matrix at different energies in a nontrivial way.
VIII Acknowledgments
We acknowledge interesting and fruitful discussions with Björn Trauzettel and Patrik Recher. This work has been funded by the German Federal Ministry of Education and Research within the funding program Photonic Research Germany (contract No. 13N14891) and by the Bavarian Ministry of Economic Affairs, Regional Development and Energy within Bavaria’s High-Tech Agenda Project “Bausteine für das Quantencomputing auf Basis topologischer Materialien mit experimentellen und theoretischen Ansätzen” (grant No. 07 02/686 58/1/21 1/22 2/23).
In Fig. 2 the single channel ABS spectrum is calculated with different sets of approximations applied, specifically for the L-junction geometry presented in Sec. VI. The regime is such that dwelling is negligible, i.e. the superconductors, as depicted in Fig. 7, are situated right at the scattering center, ${j_{\alpha}=0}$. Additionally, only the first channel is occupied as a planar mode ${n_{\mathrm{P}}=1}$ (with ${N=5}$ transversal lattice points), exhibiting strong finite dispersion, such that the chemical potential $\mu$ is tuned very closely above the minimum of that channel  (${\mu_{1}/\Delta=-1.01}$, ${\mu_{2}/\Delta=7319.5}$), as illustrated in Fig. 13.
Second, in order to compute the green curve in Fig. 2, the energy dependence of the scattering matrix (arising from finite dispersion, not dwelling) is included, but the reflection matrix is still determined within the Andreev approximation. Now, the relation (21) is no longer applicable, such that the Beenakker Equation, Eq. (18), has to be solved directly, fow which both $R$ and $S$ need to be computed first. The reflection matrix is determined by Eqs. (11) and (12). However, the scattering matrix ${S\left[E\right]}$ is now computed for the whole subgap energy range $E\in\left[-\Delta,\Delta\right]$, again numerically by Eq. (88).
Finally, the blue curve is obtained by additionally computing the reflection matrix $R$ beyond the Andreev approximation. For this, the $R$-matrix takes on a generalized form where normal backscattering submatrices are added on its diagonal, c.f. Eq. (57), and the reflection coefficients $\alpha$ no longer take on the form (12). The reflection matrix can, then, either be solved numerically by a procedure similar to the scattering problem, but by projecting onto the SN-interface instead, Eq. (92). Or, by applying the analytic expressions we where able to find for the Andreev- and backscattering coefficients in Sec. IV, Eqs. (62) and (63). For Fig. 2 we opted for the numeric version.
Figure 13: Finite dispersion in the first planar mode. Only the first channel is occupied as planar mode (solid green curve), where the chemical potential $\mu$ (green dashed line) is tuned close to the minimum of the channel, such that dispersion is non-linear for this mode. Note that here, for illustrative purposes, the green dotted line is not drawn to scale (the distances from the chemical potential $\mu$ to the minima of the first two modes actually are ${\mu_{1}/\Delta=-1.01}$ and ${\mu_{2}/\Delta=7319.5}$, repectively).
Appendix B Matching conditions at the SN-interface for the discrete model
The superconductor regions are subject to the Bogoliubov-de Gennes Hamiltonian
@@ Appendix B Matching conditions at the SN-interface for the discrete model
where $H_{N}$ is the single particle normal conductor Hamiltonian such that $\Delta=0$ for $j>j_{\alpha}$. In the discrete limit, $H_{N}$ is defined by Eq. (81) and the eigenenergies $\mathcal{E}_{k,n}$ are still in accordance with Eq. (9) but with ${\xi_{k,n}}$ defined by Eq. (83). Consequently, for subgap energies ${|E|<\Delta}$ we have
@@
where $r_{n}$ is the coefficient for normal electron-electron backscattering and $r_{\alpha,n}$ is the electron-hole Andreev reflection coefficient. As discussed in the remarks following Eq. (86), the hole reflected in positive $j$-direction still has negative momentum $-k_{\mathrm{h},n}$, since it is treated as negative charge with propagation in opposite direction. The case of an incident hole works completely analogously. Likewise, for evanescent Andreev reflection, the momenta have to be replaced as ${k\to\pm K}$ where the sign has to be chosen such that reflected components always drop off by which incoming modes always pierce the interface with the tail of the exponential, as discussed above after Eq. (86) and depicted in Figure 6.

### v2

@@ Appendix C Matching conditions at the scattering centre for the discrete model
By the above numerical procedure, the scattering matrix needs to be evaluated for each energy $E$ separately. However, we can significantly reduce computation time by exploiting the square root behaviour deduced in Sec. III.4. Thus, the scattering problem only needs to be solved for the energies ${E=0}$ and ${E=\mu_{\text{crit}}}$, and the region in between, then, can be interpolated by the square root law. Consequently, the coefficients of the electron scattering matrix can be inferred as a function of energy by the following analytic relation
@@
explicitly capture the deviation across the whole energy range from ${E=0}$ to $\mu_{\text{crit}}$. This holds in complete analogy for ${\mathrm{Arg}\left[S^{i,j}_{\mathrm{e}}\left[E\right]\right]}$, such that the full electron scattering matrix can be assembled as ${S^{i,j}_{\mathrm{e}}\left[E\right]=\left|S^{i,j}_{\mathrm{e}}\left[E\right]\right|\cdot\mathrm{exp}\left[\mathrm{Arg}\left[S^{i,j}_{\mathrm{e}}\left[E\right]\right]\right]}$. The hole components of the scattering matrix can be calculated by using particle-hole symmetry. In Fig. 9, we explicitly illustrate the very good agreement between the square root behaviour and the numerical calculation, with the example of the planar-planar reflection coefficient $R_{1,1}$. Here, the regime is such that the first channel is planar $\mu_{1}/\Delta\ll 0$, the second channel is evanescent with finite dispersion $\mu_{2}/\Delta\gtrapprox 0$, and the width of the arms is ${N=5}$ lattice sites.
Appendix D Matching conditions at the SN-interface for the discrete model
The reflection processes at the SN-interfaces can, in the discrete model, be solved by a projection procedure similar to that of the normal scattering above. However, we now need to additionally consider the superconductor regions which are subject to the Bogoliubov-de Gennes Hamiltonian
@@ Appendix D Matching conditions at the SN-interface for the discrete model
where $H_{N}$ is the single particle normal conductor Hamiltonian such that $\Delta=0$ for $j>j_{\alpha}$. In the discrete limit, $H_{N}$ is defined by Eqs. (67) and (83), and the eigenenergies $\mathcal{E}_{k,n}$ are still in accordance with Eq. (9) but with ${\xi_{k,n}}$ defined by Eq. (84). Consequently, for subgap energies ${|E|<\Delta}$ we have
@@
where $r_{n}$ is the coefficient for normal electron-electron backscattering and $r_{\alpha,n}$ is the electron-hole Andreev reflection coefficient. As discussed in the remarks following Eq. (87), the hole reflected in positive $j$-direction still has negative momentum $-k_{\mathrm{h},n}$, since it is treated as negative charge with propagation in opposite direction. The case of an incident hole works completely analogously. Likewise, for evanescent Andreev reflection, the momenta have to be replaced as ${k\to\pm\kappa}$ where the sign has to be chosen such that reflected components always drop off by which incoming modes always pierce the interface with the tail of the exponential, as discussed above after Eq. (87) and depicted in Figure 6. Projecting onto the SN-interfaces of each contact at ${j_{\mu}=j_{\alpha}}$, see also Fig. 7, yields the following set of $2N$ conditions
which correspond to the differentiability constraints in the continuum limit. Depending on which side of the interface is involved, either the ansatz for the superconductor $|\psi^{\mathrm{SC}}_{n}\rangle$ or the normal conductor region $|\psi_{n}^{\mathrm{NC}}\rangle$ is applied, accordingly. With the SN-interface being situated a finite distance of lattice points $j_{\alpha}$ from the scattering region, the wave function must be propagated in between by an operator $P$. In the absence of impurities (our default assumption for simplicity), this propagator $P$ consists of a diagonal matrix with elements ${P_{n,n}\left[j\right]=\mathrm{e}^{-K_{\nu,n}j}}$, where the momentum ${K_{\nu,n}}$ is determined as defined by Eq. (51). For instance, if $\psi^{\mathrm{in}}$ is the incident wave vector right at ${j=0}$, it would be subject to the eigenvector equation $PRP^{-1}S\psi^{\mathrm{in}}=\psi^{\mathrm{in}}$. In the standard Beenakker formalism as well as in our generalised framework above, the propagator is usually absorbed into the $S$-matrix (or, alternatively, in $R$).
While it is possible to derive a closed expression for the coefficients of the discrete SN interface, the respective formulas are impractically bulky (which is why they are not explicitly derived here, and instead the lattice model coefficients are determined numerically). For comparison and in order to simplify the calculation, we can, however, use the analytic Andreev reflection- and backscattering coefficients $\alpha_{\nu,n}$, $\beta_{\nu,n}$ from the continuous model presented in Sec. IV, cf. Eqs. (48) and (49), and simply insert the respective dispersion relations for electrons and holes from the lattice. It turns out that even if taking a conductor arm width of just ${N=5}$ lattice points the analytic approximation (solid curves) and the numerical computation (dashed curves) are in very good agreement (see Fig. 8). This is due to the fact that the chemical potential is chosen close to a change of the channel number, where the dispersion relation of the relevant channel (here the second channel, $n=2$) is to a very good accuracy parabolic.